OpenStudy (anonymous):

Help in integral: x-3 + 6/(3x+2) dx. The author's answer is x^2/2-3x + 2ln|3x+2|+C. I understand that the integral of x-3 is x^2/2-3x. But why is the integral of 6/(3x+2) equal to : 2ln|3x+2|, where does the ln comes from? thanks for your help.

OpenStudy (anonymous):

the ln come from the 1/(3x+2) you substitute 3x+2 for u u= 3x+2 du = 3 dx this changes the integral to 1/u* du/3

OpenStudy (anonymous):

6/3 = 2 multiplied by the integral of 1/u which is ln|u| you substitute back to give 2 ln|3x+2|

OpenStudy (anonymous):

oops! + C

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

yup